A 3.0 kg box is raised from rest a distance of 5.0 m by a vertical force of 90 N (a) Find the work done by the force. b) Find the work done by gravity (c) Find the final kinetic energy of the box. eBook
Answer #1

(a)

Work done by force,

W = F*d = 90*5

W = 450 J

(b)

Work done by gravity,

W = -mgd = -3*9.8*5

W = -147 J

(c)

Final KE = Wnet

KEf = 450 - 147

KEf = 303 J